# Pneumatic Multi-mode Silicone Actuator with Pressure, Vibration, and Cold Thermal Feedback

### v1

@@ III-B Material Selection and Molding Process
Ecoflex 00-30 (Smooth-On, Inc., Macungie, PA, USA), with Young’s modulus of 0.169 MPa, and a Shore hardness rating of 00-30, was selected for the construction of the actuator. This Ecoflex silicone provides $90\%$ elongation at break and a $100\%$ modulus of 10 psi, offering excellent flexibility and resilience [8]. These properties make it highly adaptable for the intended application.
The mold was made from acrylonitrile butadiene styrene (ABS) and designed to match the actuator’s shape. The overall mold consisted of two parts: an inner mold to form the actuator’s interior cavity and an outer mold to shape its external structure, as shown in Figure 4.
Figure 4: Mold designs. (a) Internal mold; (b) External mold; (c) Combined (internal and external) mold structure; (d) Internal view of the combined mold structure
As shown in Figure 5, a simple molding process was required to develop the proposed actuator. Ecoflex 00-30 liquid was poured inside the mold. After pouring the liquid material, it needed to cure for a specific amount of time until it solidified.
@@ IV Control Mechanism
The main control mechanism is based on pneumatic actuation. This requires precise airflow regulation, including the directed circulation of cold air through the lateral air nozzles of the actuator. Additionally, it involves the systematic inflation and deflation of the actuator’s two air chambers [9]. This systematic airflow regulation makes it possible to render thermal, pressure, and vibration feedback simultaneously.
@@
Second, the amplitude of the pressure sensation and the frequency of the vibration sensation are adjusted by regulating the inlet air in the air chambers. This system allows the actuator to provide responsive feedback by dynamically adjusting sensory stimuli. Four electrically controlled air solenoid valves (SC0526GC; Skoocom Technology Co. Ltd.) are used for this purpose: two valves facilitate airflow into the chambers, while the others enable air to be exhausted. Another electric pressure regulator (SMC ITV3050) controls the stiffness by managing the pressure within the air chambers.
@@
A key design emphasis is on the simultaneous control of different feedback types. The control system is designed to allow independent and simultaneous management of both the regulators and valves. This allows for separate control of each component, facilitating the simultaneous delivery of different feedback signals when desired.
@@ V Characterization of Actuator
This section evaluates the actuator’s characteristics. A series of measurements were taken, and the results were used to assess and render various types of haptic feedback. In all measurements, a silicone fingertip mock-up was inserted into the actuator to simulate a real finger. The performance of the pneumatic system relies on the actuator’s material properties and the accuracy of the pneumatic valves.
@@
Figure 8: Experimental setup for measuring vibration response.
@@ V-A1 Measurement Setup and Procedure
Figure 6 shows the experimental setup. A temperature sensor (MLX90614) was placed inside the actuator between the silicone fingertip mock-up and the cold air discharge point. Five air pressures, ranging from 3.42 to 6 bar, were used to generate varying temperatures with the Vortex tube. The flow rates of the input air ranged from 4.15 $m^{3}/h$ to 9.30 $m^{3}/h$.
V-A2 Result
Figure 7 displays the temperature transitions observed at different pressure levels over a 5-second measurement period. The room temperature was maintained at 26°C throughout the experiment, providing a consistent baseline. The minimum temperature of the generated cold air was recorded at 13°C with an air pressure of 6.00 bar. Lower air pressure results in a smaller temperature drop. Additionally, the amount of temperature drop was found to be nearly linearly related to the air pressure level, enabling precise control of the temperature. In most cases, the temperature reached saturation after 2 seconds.
@@ V-B1 Measurement Setup and Procedure
Figure 8 illustrates our actual measurement setup. An accelerometer (GY-61) was mounted between the upper air chamber and the silicone fingertip mock-up. A data collection unit (NI-DAQ 6009) recorded the acceleration data, while MATLAB was used to store the data on the PC. The actuator’s ability to provide vibrotactile feedback was evaluated across five frequencies, ranging from 10 Hz to 200 Hz. The accelerometer was calibrated prior to data collection, and with a sample rate of 1 kHz, acceleration measurements were recorded for 3 seconds at each frequency. A pressure regulator controlled the amplitude of each frequency at a pressure level of 5 psi.
V-B2 Result
Figures 9(a), and 9(b) show vibration signals under various conditions in the time and frequency domains, respectively. The maximum amplitude was observed at 80 Hz, with an acceleration of 1.7 g. The amplitude noticeably decreases as the frequency approaches 200 Hz, peaking at a value of 0.6 g. Our actuator also shows a high acceleration amplitude at the lower end of the frequency spectrum, particularly at 1 Hz, in contrast to existing vibrotactile actuators.
@@ V-C Force Response for Pressure Feedback
This section characterized the pressure feedback of our proposed actuator.
@@ V-C1 Measurement Setup and Procedure
Figure 10 shows the measurement setup. A force sensor (Wenzhou SF20) was placed above the actuation chamber of the actuator. This sensor has a force-sensing resolution of 0.01N. Data was recorded at pressure levels of 5 psi and 10 psi by adjusting the pressure regulator. Each pressure level was measured by adjusting the duration of the solenoid valve opening in 20 ms increments, from 0 ms to 240 ms.
Figure 11: Force measurement at different valve opening duration with two source pressure levels.
V-C2 Result
@@
Figure 11 illustrates the relationship between force magnitude and the valve opening duration. As expected, longer delays in the inlet pressure valve lead to greater force magnitudes. The trend peaks at 8 N with a valve opening delay of 200 ms at 10 psi, indicating that saturation occurs at this point. This behavior is attributed to the inherent pressure limitations of the pneumatic assembly when integrated with the end effector. Beyond the 8 N threshold, additional pressure primarily causes the flexible membrane to expand, increasing the contact area rather than the force. This effect persists up to 250 ms. It is important to note that keeping the solenoid valve open for more than 250 ms at 10 psi may damage the silicone air cell.
VI User Study 1: Frequency Discrimination
@@
We evaluated our system through three consecutive user studies. In the first evaluation, we focused on the actuator’s capability as a vibration feedback transducer. For general vibration rendering, particularly in simulating roughness through vibrotactile signals, the most commonly used actuator is the voice-coil-type vibrator. We tested the perceptual performance of our actuator against one of the most advanced voice-coil vibrators on the market (Haptuator BM; Tactile Lab).
@@
VI-A Participants
Fifteen participants (13 males, 2 females; average age 29) participated in the experiment. Seven had no experience with haptic rendering, while eight had general knowledge. None reported conditions affecting hand sensations. All participants gave their informed written consent to participate in this study. All experiments were approved by the Institutional Review Board (IRB) from the author’s institution.
@@ VI-A Participants
VI-B Stimuli and Experimental Design
The evaluation covers frequencies ranging from 10 Hz to 100 Hz. Five frequencies (10 Hz, 30 Hz, 50 Hz, 70 Hz, and 100 Hz) were rendered by the two actuators: our proposed pneumatic actuator and a voice coil-type actuator.
The vibration signals were delivered to the user’s dominant thimble. A within-subject design was employed to test whether the vibrations produced by the two actuators were perceived similarly.
VI-C Experimental Procedure
Before the experiment, participants received detailed instructions, which they repeated to confirm their understanding. The experiment consisted of two stages: a training session and a main session. During the training session, participants experienced vibrotactile feedback at five different frequencies using a voice coil actuator.
In the main session, participants received pairwise vibrotactile feedback: a specific frequency from the voice coil actuator and five random frequencies from the pneumatic actuator. After each trial, participants rated the haptic similarity between the two stimuli on a scale from 0 to 1 (with 0 indicating completely different). This process created a similarity matrix. Each frequency pair was assessed for 10 seconds, with a 2-second interval between pairs. To minimize distractions, audio disturbances were eliminated during the main session, allowing participants to focus solely on the haptic feedback.
VI-D Result and Discussion
Figure 12 displays the mean similarity scores for all combinations. The vertical axis represents the rendered frequencies using our actuator, while the horizontal axis indicates those rendered by the voice coil actuator. The mean similarity score for the correctly matched pairs (diagonal blocks) is 0.93.
These results suggest that the quality of frequency rendering by our actuator is comparable to that of the advanced voice coil actuator across the entire frequency range.
@@ VI-D Result and Discussion
VII User Study 2: Effect of Multi-Modal Feedback on Realism
@@
Two conditions were compared: one that utilized our multi-mode actuator to create the sensation of surface and texture through pressure and vibration and another that employed a conventional single voice-coil actuator, which provided only vibration feedback. By comparing the perceived realism between these two conditions, we aimed to investigate the impact of the additional pressure feedback on user-perceived realism.
Figure 14: Performance comparison between the proposed pneumatic and the voice coil actuator under three conditions; (a) coarse texture surface, (b) fine texture surface, and (c) 3D button.
VII-A Stimuli
The two virtual textured surfaces and the virtual push button were presented to the participants. For texture rendering, separate algorithms were required for the two actuators due to their distinct specifications and control schemes. During the interaction, the user’s fingertip position was tracked using an external optical tracker [44].
For our pneumatic actuator, a quasi-static pressure is first rendered during a contact. Simultaneously, roughness is simulated by high-frequency vibration. However, for this actuator, it is not quite possible to control the accelerations based on an arbitrary acceleration profile since the vibration is controlled by discretely opening or closing the valves for the air chamber. To address this, we employed a grid-based algorithm. Virtual grids are defined on the surface, and a short vibration burst is triggered as the user’s finger passes over each grid. The spatial frequency of the grid and the vibration amplitude determine the perceived roughness of the texture, with coarse and fine textures differing in grid frequency.
In comparison, the voice-coil actuator is more flexible in control as it gets an arbitrarily shaped waveform and produces a similar acceleration pattern. For a fair comparison, we adapted a state-of-the-art texture rendering algorithm from previous studies [45, 46, 47, 48] to our setup. This algorithm generates a noisy vibration pattern while the user strokes the surface. It dynamically adjusts the vibration pattern based on the user’s stroking velocity, with frequency and amplitude decreasing as velocity decreases. The roughness of the surface is determined by the frequency and amplitude of the underlying noise pattern. It is important to note that the voice-coil actuator does not provide static pressure for the sensation of contact.

### v2

@@ III-B Material Selection and Molding Process
Figure 3: Mold designs. (a) Internal mold; (b) External mold; (c) Combined (internal and external) mold structure; (d) Internal view of the combined mold structure.
@@
Ecoflex 00-30 (Smooth-On, Inc., Macungie, PA, USA), with Young’s modulus of 0.169 MPa and Shore hardness rating of 00-30, was selected for the construction of the actuator. This Ecoflex silicone provides $90\%$ elongation at break and a $100\%$ modulus of 10 psi, offering excellent flexibility and resilience [talhan2019tactile]. These properties make it highly adaptable for the intended application.
The mold was made of acrylonitrile butadiene styrene (ABS) and designed to match the shape of the actuator. The overall mold consisted of two parts: an inner mold to form the interior cavity of the actuator and an outer mold to shape its external structure, as shown in Figure 3.
As shown in Figure 4, a simple molding process was required to develop the proposed actuator. Ecoflex 00-30 liquid was poured into the mold. After the liquid material was poured, it needed to cure for a specific amount of time until it solidified.
@@ IV Control Mechanism
The main control mechanism is based on pneumatic actuation. This requires precise airflow regulation, including directed circulation of cold air through the lateral air nozzles of the actuator. Additionally, it involves systematic inflation and deflation of the two air chambers of the actuator [raza2024pneumatically]. This systematic airflow regulation enables the simultaneous rendering of thermal, pressure, and vibration feedback.
Figure 5: Control System architecture enabling multimodal tactile feedback (pressure, vibration, and cold thermal feedback); (a) PC: Sends control command; (b) Electronic Control Unit (ECU): Generates signals for pressure regulators, positive and negative valves; (c) Air Compressor: Supplies compressed air; (d) Electronic Pressure Regulators: Adjusts air pressure based on ECU input; (e) Vortex Tube: Produces cold air for cold thermal feedback; (f) Actuator: Delivers pressure, vibration, and cold thermal cues; (g) Pressure control Mechanism: Alternates valve openings (PVO/NVO) for static pressure; (h) Vibration control Mechanism: Rapid valve cycling creates oscillatory air pulses.
@@
Second, the amplitude of the pressure sensation and the frequency of the vibration sensation are adjusted by regulating the inlet air in the air chambers. This system allows the actuator to provide responsive feedback by dynamically adjusting the sensory stimuli. Four electrically controlled air solenoid valves (SC0526GC; Skoocom Technology Co. Ltd.) are used for this purpose: two valves facilitate airflow into the chambers, while the others enable air to be exhausted. Another electric pressure regulator (SMC ITV3050) controls the stiffness by managing the pressure within the air chambers.
@@
A key design emphasis is on the simultaneous control of different feedback types. The control system is designed to allow for independent or simultaneous management of both regulators and valves. This enables separate control of each component, facilitating the simultaneous delivery of different feedback signals when desired.
@@ V Characterization of Actuator
This section evaluates the actuator’s characteristics. A series of measurements was taken, and the results were used to assess and render various types of haptic feedback. A silicone fingertip mock-up was inserted into the actuator to simulate a real finger in all measurements. The performance of the pneumatic system relies on the actuator’s material properties and the accuracy of the pneumatic valves.
@@ V-A1 Measurement Setup and Procedure
Figure 6 shows the experimental setup. A temperature sensor (MLX90614) was placed inside the actuator between the silicone fingertip mockup and the cold air discharge point. Five air pressures, ranging from 3.42 to 6 bar, were used to generate varying temperatures with the vortex tube. The flow rates of the input air ranged from 4.15 m3/h to 9.30 m3/h.
V-A2 Result and Discussion
Figure 7 illustrates the real-time temperature response for 5 seconds at various input pressure levels. The ambient temperature was maintained at 26°C to ensure consistent baseline conditions throughout the experiment. At the highest pressure of 6.00 bar, the cold air temperature dropped rapidly, reaching approximately 13°C within the first 3 seconds. A similar but slightly less pronounced saturation trend was observed at 5.44 bar. This early stabilization indicates that thermal saturation occurs quickly at higher pressures. In contrast, at lower input pressures (3.42 bar to 4.78 bar), the temperature continued to decrease throughout the full 5-second period, with no clear saturation observed even after 4 seconds.
This behavior is consistent with the working principle of the vortex tube, where higher input pressure intensifies the vortex motion and enhances the energy separation between hot and cold air streams. As a result, higher pressures enable faster and more efficient cooling, leading to earlier saturation. In contrast, at lower pressures, the cooling capacity is reduced, resulting in a slower drop in temperature over time. In general, the results confirm that the cooling rate and saturation behavior are strongly pressure dependent, with higher pressures achieving a faster thermal response and lower steady-state temperatures.
Figure 7 also illustrates the temperature recovery phase, where the system passively returns to ambient temperature after pressure removal; higher cooling pressures show slower warming due to larger drops.
Figure 8: Experimental setup for measuring vibration response.
@@ V-B1 Measurement Setup and Procedure
Figure 8 illustrates our actual measurement setup. An accelerometer (GY-61) was mounted between the upper air chamber and the silicone fingertip mock-up. A data collection unit (NI-DAQ 6009) recorded the acceleration data, while MATLAB was used to store the data on the PC. The actuator’s ability to provide vibrotactile feedback was evaluated across five frequencies, ranging from $10200\text{\,}\mathrm{Hz}\text{/}$ The accelerometer was calibrated prior to data collection, with a sample rate of $1\text{\,}\mathrm{kHz}\text{/}$, and acceleration measurements were recorded for 3 seconds at each frequency. A pressure regulator controlled the amplitude of each frequency at a pressure level of 5 psi.
V-B2 Result and Discussion
[…]
These results show that the actuator operates most efficiently around $80\text{\,}\mathrm{Hz}\text{/}$ but maintains consistent, perceptible feedback throughout the full tested range. Additional tests carried out at very low frequencies, such as $1\text{\,}\mathrm{Hz}\text{/}$, also demonstrated strong acceleration output, highlighting the actuator’s ability to provide effective tactile feedback throughout a wide frequency range, from very low to high frequencies. At $200\text{\,}\mathrm{Hz}\text{/}$, the actuator generates $0.6\text{\,}\mathrm{g}$ of acceleration, well above the vertical vibration perception threshold of $0.0010.003\text{\,}\mathrm{g}$ [morioka2008thresholds], confirming its ability to provide detectable feedback even at high frequencies. This confirms that the actuator can provide reliable and perceptible vibrotactile feedback throughout its full operational range.
@@ V-B2 Result and Discussion
Figure 11: Force measurement at different valve opening durations with two source pressure levels.
@@ V-C Force Response for Pressure Feedback
This section characterizes the pressure feedback of the actuator.
@@ V-C1 Measurement Setup and Procedure
Figure 10 shows the measurement setup. A force sensor (Wenzhou SF20) was placed above the actuation chamber of the actuator. This sensor has a force-sensing resolution of $0.01\text{\,}\mathrm{N}\text{/}$. Data were recorded at pressure levels of 5 and 10  psi by adjusting the pressure regulator. Each pressure level was measured by adjusting the duration of the solenoid valve opening in $20\text{\,}\mathrm{ms}\text{/}$ increments, from $0\text{\,}\mathrm{ms}\text{/}$ to $240\text{\,}\mathrm{ms}\text{/}$.
V-C2 Result and Discussion
@@
Figure 11 illustrates the relationship between the magnitude of the force and the duration of valve opening. As expected, longer delays in the inlet pressure valve lead to greater force magnitudes. The trend peaks at $8\text{\,}\mathrm{N}\text{/}$ with a valve opening duration of $200\text{\,}\mathrm{ms}\text{/}$ at 10  psi, indicating that saturation occurs at this point. This behavior is attributed to the inherent pressure limitations of the pneumatic assembly when integrated with the end effector. Beyond the $8\text{\,}\mathrm{N}\text{/}$ threshold, the additional pressure primarily causes the flexible membrane to expand, increasing the contact area rather than the force. This effect persists up to $250\text{\,}\mathrm{ms}\text{/}$. It is important to note that keeping the solenoid valve open for more than $250\text{\,}\mathrm{ms}\text{/}$ at 10  psi may damage the silicone air cell.
The actuator maintains a constant force output as long as the air remains sealed inside the chamber. This stability depends on the prevention of air leakage and the complete closure of all solenoid valves.
Figure 11 also illustrates the force decay during the exhaust phase. The negative pressure valve releases the stored air within $3050\text{\,}\mathrm{ms}\text{/}$, leading to a rapid force drop. This rapid air release allows the actuator to return to its initial state quickly for repeated use.
TABLE I: Performance comparison of the proposed work with the state-of-the-art works. (N/A: Not Available; N/R: Not Reported).
Study Actuation Method Form Factor Simultaneous Actuation Max Output
Size Weight Pressure Vibration Thermal
Gallo [gallo2015flexible]
Electromagnet, Pneumatic,
Thermoelectric Cooler
23 × 23 × 6
($m$)
3 g No 200 mN N/A 24.6 - 28 °C
Talhan [talhan2019tactile]
Pneumatic
100 mm
length
4.5 g No 6.3 N 1 - 250 Hz N/A
Raza [raza2024pneumatically]
Pneumatic
55 × 55 × 20
($m$)
25 g Yes 8.3 N 1 – 250 Hz N/A
Hashem [hashem2021soft]
Pneumatic N/R N/R Yes 5.5 N 1 – 100 Hz N/A
Gallo [gallo2014design]
Hydraulic,
Thermoelectric Cooler
[…]
thick
N/R Yes 3 N 5 Hz 20 - 40 °C
Cai [cai2020thermairglove]
Hydraulic,
Thermoelectric Cooler
200 × 20 × 15
($m$)
N/R Yes 40 kpa N/A 28 – 40 °C
Lee [lee2021three]
Pneumatic,
Heater
48 × 35 × 18
($m$)
N/R Yes 11 N N/A
Room Temperature
– 90 °C
Jeong [jeong2024interactive]
Hydraulic,
Thermoelectric Cooler
N/R N/R Yes 12.3 N N/A 3 – 50 °C
Zhu [zhu2022soft]
Pneumatic,
Electro-resistive Heater
N/R N/R Yes 0.3 N N/A
Room Temperature
– 90 °C
Proposed Pneumatic
33 × 28 × 28
($m$)
9 g Yes 8 N 1 – 200 Hz
Room Temperature
13 °C
VI Performance Metrics Comparison
Table I presents a comparison of the proposed actuator with existing multi-modal haptic devices based on key performance metrics, including pressure output, vibration frequency range, thermal response, and support for simultaneous feedback.
The proposed actuator delivers a high static force of $8\text{\,}\mathrm{N}\text{/}$ while keeping a compact and lightweight form factor $33\times 28\times 28~$\mathrm{mm}\text{/}$\$, $9\text{\,}\mathrm{g}\text{/}$). This performance is comparable to that of other pneumatic systems, such as Hashem et al. [hashem2021soft] (5.5 $5\text{\,}\mathrm{N}\text{/}$) and Talhan et al. [talhan2019tactile] ($6.3\text{\,}\mathrm{N}\text{/}$). It also matches the output of Raza et al. [raza2024pneumatically], which achieved a force of $8.3\text{\,}\mathrm{N}\text{/}$. In contrast, Gallo et al. [gallo2014design] ($200\text{\,}\mathrm{mN}\text{/}$) and Zhu et al. [zhu2022soft] ($0.3\text{\,}\mathrm{N}\text{/}$) deliver a much lower force, limiting their effectiveness for fingertip feedback.
The actuator supports a wide vibration bandwidth of $1200\text{\,}\mathrm{Hz}\text{/}$, similar to Talhan [talhan2019tactile] and Raza [raza2024pneumatically], while other systems such as Cai [cai2020thermairglove] and Jeong [jeong2024interactive] do not include vibration. This allows the device to simulate a range of tactile textures.
For thermal feedback, the actuator uses a vortex cooling mechanism that rapidly lowers the contact surface temperature to 13°C in 3 seconds. While Lee et al. [lee2021three] and Zhu et al. [zhu2022soft] provide thermal feedback through heating only. On the other hand, Jeong et al. [jeong2024interactive] support both heating and cooling, but depend on external pumps and water tanks, which increases system complexity and bulk. In comparison, the proposed pneumatic actuator enables fast and effective cooling without the need for electrical heaters or fluids, offering a simpler and more compact solution for integration into wearable systems.
A key advantage is the actuator’s ability to provide pressure, vibration, and thermal feedback simultaneously using only compressed air. This reduces hardware complexity compared to systems that rely on electrical or fluid-based components, making it more suitable for immersive and responsive haptic feedback in virtual and augmented reality applications.
VII User Evaluations
To evaluate the performance and perceptual effectiveness of the proposed actuator, we conducted three user studies. The first study assessed the ability of the actuator to render different vibration frequencies. The second examined the impact of combined pressure and vibration feedback on perceived realism in surface interactions. The third explored the effectiveness of simultaneous thermo-tactile feedback in enhancing immersion within virtual scenarios.
In all user studies, participants gave their informed written consent to participate. None reported conditions affecting hand sensations. The experiments were approved by the Institutional Review Board (IRB) of the author’s institution.
VII-A User Study 1: Frequency Discrimination
@@
The first experiment focused on the actuator’s capability as a vibration feedback transducer. For general vibration rendering, particularly for simulating roughness through vibrotactile signals, the most commonly used actuator is the voice coil-type vibrator. We tested the perceptual performance of participants in identifying vibrotactile signals generated by our actuator versus a leading voice coil device (Haptuator BM; Tactile Lab).
@@
VII-A1 Participants
Fifteen participants (13 males, 2 females; average age 29) participated in the experiment. Seven had no experience with haptic rendering, while eight had general knowledge.
VII-A2 Stimuli and Experimental Design
The evaluation covers frequencies ranging from $10\text{\,}\mathrm{Hz}\text{/}$ to $100\text{\,}\mathrm{Hz}\text{/}$. Five frequencies ($10\text{\,}\mathrm{Hz}\text{/}$, $30\text{\,}\mathrm{Hz}\text{/}$, $50\text{\,}\mathrm{Hz}\text{/}$, $80\text{\,}\mathrm{Hz}\text{/}$, and $100\text{\,}\mathrm{Hz}\text{/}$) were rendered by the two actuators: our proposed pneumatic actuator and a voice coil-type actuator. Vibration signals were delivered to the dominant index finger of the user. A within-subject design was employed to test whether the vibrations produced by the two actuators were perceived similarly.
@@ VII-A2 Stimuli and Experimental Design
VII-A3 Experimental Procedure
Before the experiment, the participants received detailed instructions that they repeated to confirm their understanding. The experiment consisted of two stages: a training session and a main session. During the training session, the participants experienced vibrotactile feedback at five different frequencies using a voice coil actuator.
In the main session, participants received pairwise vibrotactile feedback: a specific frequency from the voice coil actuator and five random frequencies from the pneumatic actuator. After each trial, participants rated the haptic similarity between the two stimuli on a scale from 0 to 1 (with 0 indicating completely different). This process created a similarity matrix. Each frequency pair was assessed for 10 seconds, with a 2-second interval between pairs. Audio disturbances were eliminated during the main session to minimize distractions, allowing participants to focus solely on the haptic feedback.
VII-A4 Result and Discussion
Figure 12 displays the mean similarity scores for all combinations. The vertical axis represents the frequencies rendered using our actuator, while the horizontal axis indicates those rendered by the voice coil actuator. The mean similarity score for the correctly matched pairs (diagonal blocks) is 0.93. These results suggest that the quality of frequency rendering by our actuator is comparable to that of the advanced voice coil actuator across the entire frequency range.
@@ VII-A4 Result and Discussion
VII-B User Study 2: Effect of Multi-Modal Feedback on Realism
@@
Two conditions were compared: one that utilized our multi-mode actuator to create the sensation of surface and texture through pressure and vibration, and another that employed a conventional single voice-coil actuator, which provided only vibration feedback. By comparing perceived realism between these two conditions, we aimed to investigate the impact of additional pressure feedback on user perception of realism.
Figure 14: Performance comparison between the proposed pneumatic and the voice coil actuator under three conditions: (a) coarse texture surface, (b) fine texture surface, and (c) 3D button.
VII-B1 Stimuli
The two virtual surfaces and the push button were presented to the participants. For texture rendering, separate algorithms were required for the two actuators due to distinct specifications and control schemes. During interaction, the position of the user’s fingertip was tracked using an external optical tracker [OptiTrackV120].
For our pneumatic actuator, a quasi-static pressure is first rendered during contact. Simultaneously, roughness is simulated by high-frequency vibration. However, for this actuator, it is not quite possible to control the accelerations based on an arbitrary acceleration profile since the vibration is controlled by discretely opening or closing the valves for the air chamber. To address this, we employed a grid-based algorithm. Virtual grids are defined on the surface, and a short vibration burst is triggered as the user’s finger passes over each grid. The spatial frequency of the grid and the vibration amplitude determine the perceived roughness of the texture, with coarse and fine textures differing in grid frequency.
In comparison, the voice-coil actuator is more flexible in control as it receives an arbitrary waveform and produces a similar acceleration pattern. For a fair comparison, we adapted a state-of-the-art texture rendering algorithm from previous studies [culbertson2014modeling, abdulali2018data, nai2021vibrotactile, heravi2024development] to our setup. This algorithm generates a noisy vibration pattern while the user strokes the surface and dynamically adjusts the vibration based on stroking velocity, with frequency and amplitude decreasing as velocity decreases. The surface roughness is determined by the frequency and amplitude of the underlying noise pattern. It is important to note that the voice-coil actuator does not provide static pressure for the sensation of contact.